OpenStudy (anonymous):

Express the radical in simplest form: radical 45x

OpenStudy (ash2326):

we have \[\sqrt{45x}\] \[\sqrt{9\times 5x}\] we know \(3^2=9\) so \[\sqrt{3^2\times 5x}\] we get \[3\sqrt{5x}\]

OpenStudy (anonymous):

3√5x